OpenStudy (anonymous):

A central angle measuring 120° intercepts an arc in a circle whose radius is 3. What is length of the arc of the circle formed by this central angle? Round the length of the arc to the nearest hundredth of unit. 6.28 units 12.57 units 9.42 square units 6.28 degrees

OpenStudy (anonymous):

42 Square units. :)

OpenStudy (campbell_st):

you are after arc length... not area

OpenStudy (anonymous):

Lol thank you , but that choice isnt up there.

OpenStudy (campbell_st):

what fraction of the circle does an angle of 120 form..?

OpenStudy (anonymous):

wait what?

OpenStudy (campbell_st):

so an angle of 120 if what fraction of the circle... or revolution...?

OpenStudy (anonymous):

idk

OpenStudy (campbell_st):

ok... here is what you need  a circle is 360 degrees so \[Arc ~Length = \frac{\theta}{360} \times 2 \pi r\]